Cary is 4 years older than Dan. In 10 years the sum of their ages will be 84. Find the age of each man now.

Answer by stanbon(75887)

Cary is 4 years older than Dan. In 10 years the sum of their ages will be 84. Find the age of each man now.
-------
c = d + 4
c+10 + d+10 = 84
------
Substitute for "c" and solve for "d"::
d+4+10 + d+10 = 84
---------
2d + 24 = 84
-----
2d = 60
Dan's age = 30 yrs
-----
Cary's age = d+4 = 34 yrs
----------------
Cheers,
Stan H.
--------------